How many O atoms are in 2.00 moles of Na 2O 2?

Answers:

6.02 × 1024

2.41 × 1023

2.41 × 1024

1.20 × 1024
Answer #1

1 Na2O2 molecule = 2 O atoms

Thus, 1 mole Na2O2 molecules=2 moles O atom

Thus, 2 mole Na2O2 molecules=2*2=4 moles O atom

1 mole = 6.023*1023 atoms.

Thus, 4 moles of O atoms=4*6.023*1023 atoms=2.41*1024 atoms.(option 3) [ANSWER}